OpenStudy (anonymous):

A ball dropped from a certain height rises after bouncing to 45% of its original height. After bouncing if the ball rises to 3 ft. What is the height from which it was dropped?

OpenStudy (anonymous):

hi are you there?

OpenStudy (anonymous):

if h is the height from where it was dropped, then 45% of h=3 or .45*h=3 h=3/ .45 calculator......h=20/3 =6.667